# Active Learning Methods for Efficient Data Utilization and Model Performance Enhancement thanks:

### v1

@@ Abstract
Index Terms:
active learning, domain adaptation, multimodal learning, natural language processing, machine learning, artificial intelligence, large language models, computer vision, uncertainty sampling, query strategy, benchmarking, interpretability.
{forest}
for tree= grow=east, draw, rounded corners, minimum height=1.5em, anchor=base west, child anchor=west, parent anchor=east, base=left, rectangle, align=left, edge+=gray, line width=1pt, inner sep=2pt, font=, , where level=1fill=green!20, where level=2fill=cyan!10, where level=3fill=purple!10, where level=4fill=orange!20, where level=5fill=gray!10, [Paper Structure [Conclusion] [Applications [Specialized Tasks] [Domain Adaptation] [Natural Language Processing] [Computer Vision] ] [Evaluations [Metrics and Benchmarks] ] [Challenges [Class Imbalance] [Fairness and Bias] ] [Methodologies [Deep Learning-based Methods] [Uncertainty Estimation] ] [Background] ]
@@ I Introduction
Active Learning (AL) is a key technique in contemporary machine learning. It minimizes labeling costs while maintaining model quality or enhancing model accuracy [1, 2]. Contrary to conventional supervised learning where there are randomly labeled examples used, AL selects the most informative unlabeled instances for labeling. This improves the training and makes it scalable [3, 4]. AL is particularly useful in domains where obtaining labeled examples is difficult or costly, such as medical imaging, scientific simulations, and low-resource languages.
@@
Interest in AL has increased with the rise of deep learning. Many new techniques are being created for Deep Neural Networks (DNN) [3, 5, 6]. The techniques employ uncertainty, differences in features, and learned representations for selecting useful samples. Research also indicates that improved benchmarks and clear communication are necessary to see when AL succeeds or fails [1, 4]. Simulations demonstrate that the performance of AL will vary significantly with task, model, and data [1, 4].
@@
In principle, researchers have applied models such as Bayesian approaches [7] and Markov decision processes [8] to describe how AL chooses data over time [9]. New metrics nowadays enable comparison of the strategies adopted by AL in addition to accuracy. Examples include uncertainty, calibration, and robustness measures [4, 10]. Using these tools, the construction of more trustworthy and interpretable AL systems becomes possible.
@@
AL focus has transitioned. Previous methodologies predominantly evaluated the model. Today, more attention is on the data. Model-centered methodologies leverage model uncertainty or confidence for data selection [3, 11]. Data-centered methodologies consider the structure and diversity of the data. They employ techniques such as pseudolabeling, curriculum learning, and synthetically generated data [12, 13]. The two perspectives in combination provide a more complete description of learning between model and data.
@@
Despite advances, there are still challenges. One of them is that AL results do not necessarily replicate between tasks. Benchmarks may vary in how they normalize data, define label boundaries, or calculate measures. To address that, the community requires common standards and open-source tools [6, 10]. Another challenge is integrating AL into continual learning, where tasks and data vary over time. AL is also applied in sophisticated fields such as quantum computing and physics-based learning, where the data is expensive but of high quality [2, 14]. The uses across fields demonstrate that science will be facilitated by AL. Some of the methods of AL also model human-like behavior such as attention, curiosity, and self-regulation [15, 16]. All these concepts serve towards making more intelligent and more versatile learning systems.
@@ II Methodologies
Uncertainty Estimation and Sampling Strategies: estimation of uncertainty and sampling strategies are key aspects of AL, which facilitate the identification of informative examples for model improvement with minimal labeling. New uncertainty estimation techniques have been proposed in recent works, with evaluation of the effectiveness of prior techniques and derivation of the theoretical underpinnings of uncertainty-based AL [17]. For example, epistemic and aleatoric uncertainty are distinguished within AL, with the argument that using epistemic uncertainty for selecting informative examples is a better characterization [18]. An examination of model mismatch has shown that uncertainty-based AL underperforms random sampling when model capacity is limited [19]. Different acquisition functions have been proposed for tackling the problem. Further, non-asymptotic convergence of uncertainty sampling has been proved for binary and multi-class classification for noise-free and noisy cases [20].
@@
Deep Bayesian methods for AL have shown stronger performance on NLP tasks [21]. Siddhant and Lipton [22] demonstrate that Bayesian AL by disagreement, using uncertainty estimates from Dropout or Bayes‑by‑Backprop [23], surpasses uncertainty sampling and i.i.d baselines [24]. A belief-function‑based uncertainty method of sampling [25] reduces computational effort and captures label uncertainty by design. The advances reflect increasing attention on Bayesian and evidential approaches for active learning. All researches stress the necessity of uncertainty estimation for active learning. They investigate uncertainty sampling, Bayesian active learning, and evidential uncertainty sampling. They also mention the effect of model capacity and mismatch on performance.
@@
Convergence analysis and bounds have gained prominent focus. Lu et al. [26] re-benchmarked pool-based active learning for binary classification. Mussmann and Liang [27] explored the relationship between error and data efficiency for uncertainty sampling. Other research suggests other approaches like weighted uncertainty sampling [28] and learning the model uncertainty [29]. A stopping criteria using deterministic generalization bounds has been proposed to automatically stop active learning [30]. All these contrasting opinions reflect the controversy in the field.
@@
However, model capacity and mismatch are still important challenges [19]. Computational intensity is also an issue. Simplifications of the computational task have become available [25]. There are limited theoretical guarantees as well. There remains a need for more work on the principles and limitations of active learning.
@@
In summary, recent research has significantly improved our knowledge of uncertainty estimation and sampling techniques in AL. The creation of novel uncertainty estimation techniques, model mismatch analysis, and the provision of theoretical bounds have all improved our understanding in this regard. Nonetheless, ongoing challenges and limitations serve as evidence that there remains room for further research in AL studies, specifically towards addressing model capacity complexities, computational complexities, and theoretical underpinnings. As research in AL proceeds, it will likely birth new advances, drawing from the work of these recent studies [25, 19, 20].
@@
TABLE I: Comparison between BALD and uniform labeling after 10% of the dataset has been labeled. $\lambda$ is the weight of the gradient reversal term. Both metrics are evaluated on a balanced held-out set. Standard deviation is reported by repeating the experiment with 3 different random seeds [32].
@@
Deep Learning-based Active Learning Methods: deep learning-based AL garnered considerable attention over the past years [33, 34, 35, 36, 37, 38]. However, numerous approaches do not work and are even inferior to random sampling. Entropy-based AL remains a solid baseline and usually outperforms single-model methods, and the pretraining options, initial budget, and budget step size can have unexpected effects on results [39].
@@
Some research uses cutting-plane algorithms for deep networks. These are a decent substitute for gradient-based alternatives [40]. AL also employs pseudo-labels over real labels. Incremental and cumulative learning assist with enhanced performance [3]. Evaluation of AL for real-world tasks such as semi-supervised learning and object detection is crucially important [39, 41]. Li et al. also identify that we still have no comprehensive survey that explains deep AL methodologies, their benefits and drawbacks, as well as applications [42].
@@
Deep AL is extremely sensitive to the training conditions, like data augmentation [41]. It can enhance performance if coupled with semi-supervised learning or gradient-free optimization [40, 3]. A study reveals model-centric (model-centered and query-strategy-centered) and data-centric (pseudo-labels and training modes-centered) aspects of active learning with deep neural network models, e.g., training mode, query strategies, availability of unlabeled data, initial training, and network specifications [3]. Despite the sophisticated alternatives, there remains controversy regarding whether simple entropic approaches outperform alternatives [39, 42].
@@
Technical aspects also play a part. The application of ReLU networks and network depth influences AL performance in image classification and segmentation tasks [40]. It’s also necessary to test approaches on both real and synthetic datasets in order to see what their limits are [40, 3]. Stating the statistical properties of chosen examples and test error estimation also serves to estimate deep AL [3].
@@
There has been improvement, but there are still challenges. We have not yet had a full comparison of various deep AL techniques [42]. The outcomes depend heavily on training and experiments [41]. There exists a great demand for AL techniques that perform effectively with big datasets and few labels [39, 40]. Even more, approaching the design of AL techniques that achieve the optimal solution reliably remains a major challenge [40].
@@
Interactive and Question-Driven Learning in Active Learning: this is an important part of active learning. It helps machines ask questions and learn through dialogue. For example, the INTERACT framework lets large language models join in student-teacher conversations. This leads to up to 25% better performance [43]. It shows that interaction is valuable—machines learn better when they ask and receive information step by step. Another method, structural query-by-committee [44], gives a general algorithm for interactive learning tasks. It works well even when there is noise in the data.
@@
The difference between passive and active learning is a common topic. Active learning usually gives better results [43, 15]. Some tools compare different strategies, helping users see their pros and cons [45]. Other studies compare human-like learning with machine-based systems like INTERACT [43, 15].
@@
Many papers show that human-like active learning, interactive, question-based learning performs better than passive learning in both human and machine studies [15]. This kind of learning also focuses on explainability. Visualization tools help people understand how models learn and make decisions [45]. Some papers also connect active learning with curiosity. They suggest that active learning is like curiosity-driven exploration [46, 15].
@@
On the technical side, several methods are used. These include student-teacher dialogues [43], query-by-committee [44], and knowledge distillation [15]. Tools like the visual interface in [45] give useful views into how training happens. Still, there are problems. Scalability is one issue [43]. Dealing with noise and uncertainty is another [44]. Also, better evaluation metrics are needed [45]. Making humans and machines work together is also a challenge [15].
@@
Interactive active learning can be applied in natural language processing [5], visual question answering [47], and education [48]. But to make more progress, we need to fix its current problems, e.g., building better metrics and connecting human-style and machine-style learning. In the future, interactive learning may change how machines learn—making them faster, more accurate, and easier to understand [49, 9, 50].
@@ III Challenges
Fairness and Bias in Active Learning: recent research indicates that careless collection of data will exacerbate prevalent biases [51]. Label bias in data collection will be augmented by bias despite fairness constraints [51]. Uncertainty-based heuristics such as Bayesian Active Learning by Disagreement (BALD) enhance predictive parity and accuracy over i.i.d. sampling [32]. New weights can correct and formalize statistical bias in AL [52].
Fair learning frameworks like Falcon employ multi-armed bandits for selecting informative examples and obtaining state-of-the-art performance [53]. Bayesian adaptive experimental design has the disadvantage of model misspecification bias, but the inclusion of noise in the model alleviates the problem [54]. Novel weights correct statistical bias [52]. Bayesian adaptive experimental design is evaluated for its robustness to bias, and experiments test the effectiveness of fair AL approaches [54]. All works emphasize controlling label bias, statistical bias, and algorithmic bias in AL.
While there has been progress, challenges persist. It is challenging to deal with fairness and bias in AL and must account for several sources of bias [51]. Misspecification and noise impede the robustness of Bayesian active learning design [54]. There is a trade-off between informativeness and postpone rate according to Falcon [53]. Scalability and efficacy are also crucial for fair AL methods [55, 56]. Overall, the development of fair active learning approaches has significant implications for a wide range of applications, e.g., natural language processing [55] and educational AI [57].
Figure 4: Illustration of active learning procedure implemented by ALBench [58] based on YMIR [59]. Green indicates labeled data, yellow indicates mined data, and red indicates raw data.
Figure 5: Desai et al. [60] proposed framework integrating weak supervision into the active learning pipeline. It features an adaptive supervision module that dynamically escalates the level of supervision, enabling the transition to stronger supervision modes when necessary during model training.
Class Balance and Imbalance Issues in Active Learning: class imbalance hurts classifier performance in AL. It degrades the minority classes. Recent research pointed out that it is a potential factor in producing suboptimal classifiers and poor minority class performance [61, 62]. Researchers propose methods for handling class imbalance, including optimization frameworks that have a class-balancing feature [61], illustrated below:
Algorithm 1 Class Balancing AL [61]
@@
Algorithm 2 Greedy Class Balancing Selection [61]
@@
[…]
@@
where $\lambda$ is a parameter that regularizes the contribution of the balancing term in the objective. The cost function consists of an affine term and an $\ell_{1}$ norm that are both convex, and subsequently, their linear combination is also convex. However, as the constraint is non-convex, the optimization problem becomes non-convex. The underlying problem is Binary Programming [63] that can be optimally solved by an off-the-shelf optimizer using LP relaxation [64] and the branch-and-bound method.
@@
Algorithm 2 (Greedy Class Balancing Selection) provides an efficient, iterative approach to sample selection. At each step, the algorithm identifies the most informative sample by balancing two criteria: selecting samples far from existing labeled samples (based on a distance metric $D$), and maintaining class balance (guided by threshold $\Omega(c)$). Specifically, the algorithm greedily minimizes an objective that combines these two criteria. After each selection, the algorithm updates the distance matrix and continues until the budget per cycle is reached. In addition, others also include algorithms that spot class separation thresholds and pick the most uncertain instances [62], and graph-based techniques merge graph-based AL and deep learning [65].
@@
Studies highlight the necessity of addressing various types of imbalance, such as proportion, variance, distance, neighborhood, and quality imbalance between/among classes [66]. They enumerate types of imbalance: proportion, variance, distance, neighborhood, and quality. Real-world problems frequently encounter such imbalances [67, 68]. Current AL techniques may not perform or may deteriorate on imbalanced data, highlighting the necessity for new techniques and approaches [65, 66].
@@
Graph and deep learning methods are also under investigation as possible solutions for class imbalance in AL, with research showing their value in such applications as image classification [61] and NLP tasks [67]. The method of choice will vary with task and data [69, 70]. Researchers are also investigating a variety of technical solutions, such as optimization frameworks [61], new algorithms based on separation thresholds [62], and meta-learning for learning hyperparameters pertaining to imbalance [66].
@@
Certain works propose new algorithms for treating imbalanced AL [62, 65]. Others provide theoretical suggestions and promote a rethink of class imbalance [66]. There are different opinions regarding the necessity of class balance, with some believing that it’s crucial [65], and others proposing that there are ways of handling class imbalance by other means [62].
@@
Finally, effective solutions for managing class imbalance in AL will depend on further insight into the intricate relationships between learning strategies, class balance, and characteristics of the data. Effective methods necessitate further insight into how learning strategies, class balance, and data characteristics interplay. Uncovering these relationships will provide stronger systems for real-world tasks [71, 72]. As technology advances, we anticipate improved methods for coping with class imbalance. Better methods will enhance performance on applications.

### v2

@@ Abstract
[…]
Paper Structure
Conclusion
Applications
Specialized Tasks
Domain Adaptation
Natural Language Processing
Computer Vision
Evaluations
Metrics and Benchmarks
Challenges
Class Imbalance
Fairness and Bias
Methodologies
Deep Learning-based Methods
Uncertainty Estimation
Background
@@ I Introduction
Active Learning (AL) is a key technique in contemporary machine learning. It minimizes labeling costs while maintaining model quality or enhancing model accuracy [10.1007/978-3-642-41398-8_16, ding2023active]. Contrary to conventional supervised learning where there are randomly labeled examples used, AL selects the most informative unlabeled instances for labeling. This improves the training and makes it scalable [bosser2021model, abraham2020rebuilding]. AL is particularly useful in domains where obtaining labeled examples is difficult or costly, such as medical imaging, scientific simulations, and low-resource languages.
@@
Interest in AL has increased with the rise of deep learning. Many new techniques are being created for Deep Neural Networks (DNN) [bosser2021model, zhang2022survey, jukic-snajder-2023-smooth]. The techniques employ uncertainty, differences in features, and learned representations for selecting useful samples. Research also indicates that improved benchmarks and clear communication are necessary to see when AL succeeds or fails [10.1007/978-3-642-41398-8_16, abraham2020rebuilding]. Simulations demonstrate that the performance of AL will vary significantly with task, model, and data [10.1007/978-3-642-41398-8_16, abraham2020rebuilding].
@@
In principle, researchers have applied models such as Bayesian approaches [kruschke2008bayesian] and Markov decision processes [puterman1990markov] to describe how AL chooses data over time [2306.08001]. New metrics nowadays enable comparison of the strategies adopted by AL in addition to accuracy. Examples include uncertainty, calibration, and robustness measures [abraham2020rebuilding, jiang2024actively]. Using these tools, the construction of more trustworthy and interpretable AL systems becomes possible.
@@
AL focus has transitioned. Previous methodologies predominantly evaluated the model. Today, more attention is on the data. Model-centered methodologies leverage model uncertainty or confidence for data selection [bosser2021model, 2201.03947]. Data-centered methodologies consider the structure and diversity of the data. They employ techniques such as pseudolabeling, curriculum learning, and synthetically generated data [chu2016can, 2304.05246]. The two perspectives in combination provide a more complete description of learning between model and data.
@@
Despite advances, there are still challenges. One of them is that AL results do not necessarily replicate between tasks. Benchmarks may vary in how they normalize data, define label boundaries, or calculate measures. To address that, the community requires common standards and open-source tools [jukic-snajder-2023-smooth, jiang2024actively]. Another challenge is integrating AL into continual learning, where tasks and data vary over time. AL is also applied in sophisticated fields such as quantum computing and physics-based learning, where the data is expensive but of high quality [ding2023active, 2302.14567]. The uses across fields demonstrate that science will be facilitated by AL. Some of the methods of AL also model human-like behavior such as attention, curiosity, and self-regulation [2011.03733, takezoe2023deep]. All these concepts serve towards making more intelligent and more versatile learning systems.
@@ II Methodologies
Uncertainty Estimation and Sampling Strategies: estimation of uncertainty and sampling strategies are key aspects of AL, which facilitate the identification of informative examples for model improvement with minimal labeling. New uncertainty estimation techniques have been proposed in recent works, with evaluation of the effectiveness of prior techniques and derivation of the theoretical underpinnings of uncertainty-based AL [li2024adaptive]. For example, epistemic and aleatoric uncertainty are distinguished within AL, with the argument that using epistemic uncertainty for selecting informative examples is a better characterization [nguyen2019epistemic]. An examination of model mismatch has shown that uncertainty-based AL underperforms random sampling when model capacity is limited [2408.13690]. Different acquisition functions have been proposed for tackling the problem. Further, non-asymptotic convergence of uncertainty sampling has been proved for binary and multi-class classification for noise-free and noisy cases [raj2022convergence].
@@
Deep Bayesian methods for AL have shown stronger performance on NLP tasks [carvalho2024deep]. Siddhant and Lipton [1808.05697] demonstrate that Bayesian AL by disagreement, using uncertainty estimates from Dropout or Bayes‑by‑Backprop [hernandez2015probabilistic], surpasses uncertainty sampling and i.i.d baselines [he2021towards]. A belief-function‑based uncertainty method of sampling [hoarau2024evidential] reduces computational effort and captures label uncertainty by design. The advances reflect increasing attention on Bayesian and evidential approaches for active learning. All researches stress the necessity of uncertainty estimation for active learning. They investigate uncertainty sampling, Bayesian active learning, and evidential uncertainty sampling. They also mention the effect of model capacity and mismatch on performance.
@@
Convergence analysis and bounds have gained prominent focus. Lu et al. [2306.08954] re-benchmarked pool-based active learning for binary classification. Mussmann and Liang [mussmann2018relationship] explored the relationship between error and data efficiency for uncertainty sampling. Other research suggests other approaches like weighted uncertainty sampling [1909.04928] and learning the model uncertainty [kye2023tidal]. A stopping criteria using deterministic generalization bounds has been proposed to automatically stop active learning [2005.07402]. All these contrasting opinions reflect the controversy in the field.
@@
However, model capacity and mismatch are still important challenges [2408.13690]. Computational intensity is also an issue. Simplifications of the computational task have become available [hoarau2024evidential]. There are limited theoretical guarantees as well. There remains a need for more work on the principles and limitations of active learning.
@@
In summary, recent research has significantly improved our knowledge of uncertainty estimation and sampling techniques in AL. The creation of novel uncertainty estimation techniques, model mismatch analysis, and the provision of theoretical bounds have all improved our understanding in this regard. Nonetheless, ongoing challenges and limitations serve as evidence that there remains room for further research in AL studies, specifically towards addressing model capacity complexities, computational complexities, and theoretical underpinnings. As research in AL proceeds, it will likely birth new advances, drawing from the work of these recent studies [hoarau2024evidential, 2408.13690, raj2022convergence].
@@
TABLE I: Comparison between BALD and uniform labeling after 10% of the dataset has been labeled. $\lambda$ is the weight of the gradient reversal term. Both metrics are evaluated on a balanced held-out set. Standard deviation is reported by repeating the experiment with 3 different random seeds [2104.06879].
@@
Deep Learning-based Active Learning Methods: deep learning-based AL garnered considerable attention over the past years [lewis1994heterogeneous, lewis1995sequential, luo2013latent, settles2008analysis, vijayanarasimhan2014large, iglesias2011combining]. However, numerous approaches do not work and are even inferior to random sampling. Entropy-based AL remains a solid baseline and usually outperforms single-model methods, and the pretraining options, initial budget, and budget step size can have unexpected effects on results [2403.14800].
@@
Some research uses cutting-plane algorithms for deep networks. These are a decent substitute for gradient-based alternatives [2410.02145]. AL also employs pseudo-labels over real labels. Incremental and cumulative learning assist with enhanced performance [bosser2021model]. Evaluation of AL for real-world tasks such as semi-supervised learning and object detection is crucially important [2403.14800, 1912.05361]. Li et al. also identify that we still have no comprehensive survey that explains deep AL methodologies, their benefits and drawbacks, as well as applications [li2024survey].
@@
Deep AL is extremely sensitive to the training conditions, like data augmentation [1912.05361]. It can enhance performance if coupled with semi-supervised learning or gradient-free optimization [2410.02145, bosser2021model]. A study reveals model-centric (model-centered and query-strategy-centered) and data-centric (pseudo-labels and training modes-centered) aspects of active learning with deep neural network models, e.g., training mode, query strategies, availability of unlabeled data, initial training, and network specifications [bosser2021model]. Despite the sophisticated alternatives, there remains controversy regarding whether simple entropic approaches outperform alternatives [2403.14800, li2024survey].
@@
Technical aspects also play a part. The application of ReLU networks and network depth influences AL performance in image classification and segmentation tasks [2410.02145]. It’s also necessary to test approaches on both real and synthetic datasets in order to see what their limits are [2410.02145, bosser2021model]. Stating the statistical properties of chosen examples and test error estimation also serves to estimate deep AL [bosser2021model].
@@
There has been improvement, but there are still challenges. We have not yet had a full comparison of various deep AL techniques [li2024survey]. The outcomes depend heavily on training and experiments [1912.05361]. There exists a great demand for AL techniques that perform effectively with big datasets and few labels [2403.14800, 2410.02145]. Even more, approaching the design of AL techniques that achieve the optimal solution reliably remains a major challenge [2410.02145].
@@
Interactive and Question-Driven Learning in Active Learning: this is an important part of active learning. It helps machines ask questions and learn through dialogue. For example, the INTERACT framework lets large language models join in student-teacher conversations. This leads to up to 25% better performance [2412.11388]. It shows that interaction is valuable—machines learn better when they ask and receive information step by step. Another method, structural query-by-committee [1803.06586], gives a general algorithm for interactive learning tasks. It works well even when there is noise in the data.
@@
The difference between passive and active learning is a common topic. Active learning usually gives better results [2412.11388, 2011.03733]. Some tools compare different strategies, helping users see their pros and cons [2111.04936]. Other studies compare human-like learning with machine-based systems like INTERACT [2412.11388, 2011.03733].
@@
Many papers show that human-like active learning, interactive, question-based learning performs better than passive learning in both human and machine studies [2011.03733]. This kind of learning also focuses on explainability. Visualization tools help people understand how models learn and make decisions [2111.04936]. Some papers also connect active learning with curiosity. They suggest that active learning is like curiosity-driven exploration [1312.2936, 2011.03733].
@@
On the technical side, several methods are used. These include student-teacher dialogues [2412.11388], query-by-committee [1803.06586], and knowledge distillation [2011.03733]. Tools like the visual interface in [2111.04936] give useful views into how training happens. Still, there are problems. Scalability is one issue [2412.11388]. Dealing with noise and uncertainty is another [1803.06586]. Also, better evaluation metrics are needed [2111.04936]. Making humans and machines work together is also a challenge [2011.03733].
@@
Interactive active learning can be applied in natural language processing [zhang2022survey], visual question answering [1711.01732], and education [2406.13903]. But to make more progress, we need to fix its current problems, e.g., building better metrics and connecting human-style and machine-style learning. In the future, interactive learning may change how machines learn—making them faster, more accurate, and easier to understand [misra2018learning, 2306.08001, 2202.00254].
@@ III Challenges
Fairness and Bias in Active Learning: recent research indicates that careless collection of data will exacerbate prevalent biases [li2022more]. Label bias in data collection will be augmented by bias despite fairness constraints [li2022more]. Uncertainty-based heuristics such as Bayesian Active Learning by Disagreement (BALD) enhance predictive parity and accuracy over i.i.d. sampling [2104.06879]. New weights can correct and formalize statistical bias in AL [2101.11665].
Fair learning frameworks like Falcon employ multi-armed bandits for selecting informative examples and obtaining state-of-the-art performance [10.14778/3641204.3641207]. Bayesian adaptive experimental design has the disadvantage of model misspecification bias, but the inclusion of noise in the model alleviates the problem [2205.13698]. Novel weights correct statistical bias [2101.11665]. Bayesian adaptive experimental design is evaluated for its robustness to bias, and experiments test the effectiveness of fair AL approaches [2205.13698]. All works emphasize controlling label bias, statistical bias, and algorithmic bias in AL.
While there has been progress, challenges persist. It is challenging to deal with fairness and bias in AL and must account for several sources of bias [li2022more]. Misspecification and noise impede the robustness of Bayesian active learning design [2205.13698]. There is a trade-off between informativeness and postpone rate according to Falcon [10.14778/3641204.3641207]. Scalability and efficacy are also crucial for fair AL methods [2302.05711, 2312.08559]. Overall, the development of fair active learning approaches has significant implications for a wide range of applications, e.g., natural language processing [2302.05711] and educational AI [2407.18745].
Figure 4: Illustration of active learning procedure implemented by ALBench [2207.13339] based on YMIR [huang2021ymir]. Green indicates labeled data, yellow indicates mined data, and red indicates raw data.
Figure 5: Desai et al. [1908.02454] proposed framework integrating weak supervision into the active learning pipeline. It features an adaptive supervision module that dynamically escalates the level of supervision, enabling the transition to stronger supervision modes when necessary during model training.
Class Balance and Imbalance Issues in Active Learning: class imbalance hurts classifier performance in AL. It degrades the minority classes. Recent research pointed out that it is a potential factor in producing suboptimal classifiers and poor minority class performance [bengar2022class, 2312.09196]. Researchers propose methods for handling class imbalance, including optimization frameworks that have a class-balancing feature [bengar2022class], illustrated below:
Algorithm 1 Class Balancing AL [bengar2022class]
@@
Algorithm 2 Greedy Class Balancing Selection [bengar2022class]
@@
$\textrm{s.t.}\quad\boldsymbol{z}^{T}\mathds{1}_{N\times 1}=b,\quad z_{i}\in\{0,1\},\quad\forall i=1,2,.,N,$
@@
where $\lambda$ is a parameter that regularizes the contribution of the balancing term in the objective. The cost function consists of an affine term and an $\ell_{1}$ norm that are both convex, and subsequently, their linear combination is also convex. However, as the constraint is non-convex, the optimization problem becomes non-convex. The underlying problem is Binary Programming [soudi1998optimized] that can be optimally solved by an off-the-shelf optimizer using LP relaxation [yanover2006linear] and the branch-and-bound method.
@@
Algorithm 2 (Greedy Class Balancing Selection) provides an efficient, iterative approach to sample selection. At each step, the algorithm identifies the most informative sample by balancing two criteria: selecting samples far from existing labeled samples (based on a distance metric $D$), and maintaining class balance (guided by threshold $\Omega(c)$). Specifically, the algorithm greedily minimizes an objective that combines these two criteria. After each selection, the algorithm updates the distance matrix and continues until the budget per cycle is reached. In addition, others also include algorithms that spot class separation thresholds and pick the most uncertain instances [2312.09196], and graph-based techniques merge graph-based AL and deep learning [zhang2022galaxy].
@@
Studies highlight the necessity of addressing various types of imbalance, such as proportion, variance, distance, neighborhood, and quality imbalance between/among classes [2305.03900]. They enumerate types of imbalance: proportion, variance, distance, neighborhood, and quality. Real-world problems frequently encounter such imbalances [2210.04675, zhao2021active]. Current AL techniques may not perform or may deteriorate on imbalanced data, highlighting the necessity for new techniques and approaches [zhang2022galaxy, 2305.03900].
@@
Graph and deep learning methods are also under investigation as possible solutions for class imbalance in AL, with research showing their value in such applications as image classification [bengar2022class] and NLP tasks [2210.04675]. The method of choice will vary with task and data [choi2021vab, gilhuber2023overcome]. Researchers are also investigating a variety of technical solutions, such as optimization frameworks [bengar2022class], new algorithms based on separation thresholds [2312.09196], and meta-learning for learning hyperparameters pertaining to imbalance [2305.03900].
@@
Certain works propose new algorithms for treating imbalanced AL [2312.09196, zhang2022galaxy]. Others provide theoretical suggestions and promote a rethink of class imbalance [2305.03900]. There are different opinions regarding the necessity of class balance, with some believing that it’s crucial [zhang2022galaxy], and others proposing that there are ways of handling class imbalance by other means [2312.09196].
@@
Finally, effective solutions for managing class imbalance in AL will depend on further insight into the intricate relationships between learning strategies, class balance, and characteristics of the data. Effective methods necessitate further insight into how learning strategies, class balance, and data characteristics interplay. Uncovering these relationships will provide stronger systems for real-world tasks [1407.1124, aaboud2018search]. As technology advances, we anticipate improved methods for coping with class imbalance. Better methods will enhance performance on applications.